(1+yz)dx + x(z-x)dy -(1+xy)dz=0
1
Expert's answer
2020-10-14T15:12:30-0400

"\displaystyle\textsf{This is a Pfaffian differential equation in three}
\\\textsf{variables and we must verify its integrabilty}
\\\textsf{and determine its primitive.}\\

(1 + yz)\mathrm{d}x + (xz − x^2)\mathrm{d}y − (1 + xy)\mathrm{d}z = 0 \\

\textsf{The necessary and sufficient condition}\\\textsf{for iintegrability is} \\

\textbf{X}\cdot curl\textbf{X} = 0 \\

\textbf{X}=(1 + yz,xz − x^2,-1 - xy), \textsf{so that}\\

 \nabla \times X = \begin{vmatrix}
\textbf{i} &\textbf{j} &\textbf{k} \\
\frac{\partial}{\partial x} &\frac{\partial}{\partial y} &\frac{\partial}{\partial z}\\
1 + yz &xz − x^2 &-1 - xy
\end{vmatrix} = \\

\begin{aligned}
&\textbf{i} \left(\frac{\partial}{\partial y}(-1 - xy) - \frac{\partial}{\partial z}(xz - x^2)\right) - 
\\&\textbf{j} \left(\frac{\partial}{\partial x}(-1 - xy) - \frac{\partial}{\partial z}(1 + yz)\right) + 
\\&\textbf{k} \left(\frac{\partial}{\partial x}(xz - x^2) - \frac{\partial}{\partial y}(1 + yz)\right)
\end{aligned} \\

\begin{aligned}
&=\textbf{i}(-x - x) - \textbf{j}(-y - y) + \textbf{k}(z - 2x - z) \\&= -2x\textbf{i} + 2y\textbf{j} - 2x\textbf{k}
\end{aligned} \\

\begin{aligned}
\therefore &(1 + yz,xz − x^2,-1 -xy) \cdot (-2x, 2y, -2x) \\&= -2x(1 + yz) + 2y(xz - x^2) + 2x(1 + xy) \\&= -2x - 2xyz + 2xyz - 2yx^2 + 2x + 2x^2y = 0
\end{aligned} \\

\textsf{Thus, the given equation is integrable.}\\

\textsf{Solving by Inspection} \\

(1 + yz)\mathrm{d}x + (xz − x^2)\mathrm{d}y − (1 + xy)\mathrm{d}z = 0\\

(1 + yz)\mathrm{d}x + (xz − x^2)\mathrm{d}y − (1 + xy)\mathrm{d}z = 0\\

-(1 + yz)\mathrm{d}x - (xz − x^2)\mathrm{d}y − (1 + xy)\mathrm{d}z = 0\\

(1 + yx)\mathrm{d}z - (1 + yx)\mathrm{d}x - x(z - x) \mathrm{d}y - y(z - x)\mathrm{d}x = 0\\

(1 + yx)\mathrm{d}(z - x) - (x\mathrm{d}y + y\mathrm{d}x)(z - x) = 0\\

(1 + yx)\mathrm{d}(z - x) - \mathrm{d}(1 + xy)(z - x) = 0 \\

\implies \frac{\mathrm{d}(z - x)}{z - x} - \frac{\mathrm{d}(1 + xy)}{1 + xy} = 0\\

\textsf{Integrating both sides, we have}\\

\int\frac{\mathrm{d}(z - x)}{z - x} - \int\frac{\mathrm{d}(1 + xy)}{1 + xy} = \int0\, \mathrm{d}x\\

\ln(z - x) - \ln(1 + xy) = C_1 \\

\ln(z - x) = \ln(A(1 + xy)), \hspace{1cm} \{C_1 = \ln(A)\}\\

\implies z = x + A(1 + xy) \\



\therefore z = x + A (1 + yx)\hspace{0.1cm}\textsf{is a solution to the Pfaffian differential equation}"